solve 5y-8+y=9y+23-6y I am stuck on this question any help would be greatly appreciated.
Answer by user_dude2008(1862)

5y-8+y=9y+23-6y

6y-8=3y+23

6y-3y=23+8

3y=31

y=31/3 <--------------------------- Answer